# LLM-Guided Scenario-based GUI Testing

### v2

@@ 4.6. RQ5: Bug Detection Capability
Table 7. Result of RQ5: Bug Detection Capability
@@ 6.1. Vision-based GUI Testing
Early efforts primarily focus on visual defect detection and widget recognition. Liu et al. (Liu et al., 2020) propose OwlEye, a tool that detects display inconsistencies in GUI screenshots. White et al. (White et al., 2019) apply machine learning techniques to identify GUI widgets, thereby improving the efficiency of random testing strategies. Xie et al. (Xie et al., 2020) introduce UIED, which integrates traditional computer vision with deep learning to enhance widget detection accuracy, while Zhang et al. (Zhang et al., 2021) develop a more efficient model for GUI element detection. Mansur et al. (Hasan Mansur et al., 2023) present AidUI, which combines CV and NLP techniques to detect dark patterns, improving the ethical and usability evaluation of user interfaces. In addition, Baral et al. (Baral et al., 2024) automate the construction of GUI-based test oracles for mobile apps, further illustrating the potential of vision-driven approaches in improving testing reliability.
@@ Acknowledgements.
The authors would like to thank the editors and anonymous reviewers for their time and comments. This work is supported partially by the National Natural Science Foundation of China (62141215, 62272220, 62372228).

### v3

@@ 4.3. RQ2: Bug Detection Capability
Table 3. Result of RQ2: Bug Detection Capability
@@
4.4. RQ3: Widget Localization Effectiveness
Widget localization plays a crucial role in enabling accurate and reliable GUI interactions during automated testing. Unlike pure vision-based or text-based methods, ScenGen integrates the pixel-level precision of traditional computer vision (CV) algorithms with the semantic reasoning capabilities of multi-modal LLMs to handle visually complex and context-rich user interfaces. RQ3 aims to evaluate the effectiveness of this hybrid strategy in widget matching and localization, examining how the integration of CV techniques and LLM-driven understanding improves both accuracy and robustness in real-world testing environments.
Table 4. Result of RQ3: Widget Localization Effectiveness
Scenario N $n_{1}$ $n_{f}$ $Acc_{1}$ $Acc_{f}$ $\uparrow$
Email 33 28 32 84.85% 96.97% 14.29%
Music 22 12 22 54.55% 100.00% 83.33%
Note 53 44 52 83.02% 98.11% 18.18%
Photo 9 6 9 66.67% 100.00% 50.00%
Translation 35 34 34 97.14% 97.14% 0.00%
Shopping 31 26 30 83.87% 96.77% 15.38%
Login 35 32 33 91.43% 94.29% 3.13%
Calculation 44 34 44 77.27% 100.00% 29.41%
Weather 27 21 26 77.78% 96.30% 23.81%
Alarm-clock 23 21 23 91.30% 100.00% 9.52%
Sum/Avg 312 258 305 80.79% 97.96% 24.71%
The experimental results are summarized in Table 4. Across ten functional scenarios, the initial localization accuracy ($Acc_{1}$) ranges from 54.55% to 97.14%, while the final localization accuracy ($Acc_{f}$) after self-correction improves substantially, reaching between 94.29% and 100%. In four scenarios, the final accuracy achieves a perfect 100%. On average, the initial accuracy is 80.79%, and after correction, the average rises to 97.76%, with all scenarios over 94% in the final stage.
Furthermore, qualitative analysis reveals that most initial localization errors occur in cases where widgets are visually occluded (e.g., drop-down or scrollable menus) or contain highly similar visual components (e.g., repeated icons). The self-correction mechanism successfully resolves these issues by leveraging LLM-based semantic reasoning to re-evaluate contextual cues and by refining coordinate-level predictions based on CV feedback. This cooperative process effectively mimics human testers’ behavior when re-targeting misclicked or ambiguous GUI elements. Overall, RQ3 demonstrates that combining computer vision algorithms with multi-modal LLM reasoning significantly enhances widget localization reliability in GUI testing. The self-correction mechanism further strengthens this process by dynamically adapting to interface changes and mitigating recognition errors. These results validate that ScenGen achieves precise, context-aware localization across diverse GUI layouts, ensuring a stable foundation for accurate and scenario-driven automated test execution.
4.5. RQ4: Logical Decision-Making Effectiveness
RQ4 evaluates whether the Decider can accurately interpret the requirements of a testing scenario and generate appropriate next-step operations based on the current app GUI state. The goal is to assess both the logical soundness of initial decisions and the effectiveness of the self-correction mechanism when feedback from the environment necessitates revision. We use decision accuracy as the primary evaluation metric. To measure the impact of the self-correction mechanism, each decision and its potential subsequent correction are treated as a single decision unit. We then calculate the initial accuracy ($Acc_{1}$) and final accuracy ($Acc_{f}$) according to:
Table 5. Result of RQ4: Logical Decision-Making Effectiveness
Scenario N $n_{1}$ $n_{f}$ $Acc_{1}$ $Acc_{f}$ $\uparrow$
Email 39 38 39 97.44% 100.00% 2.63%
Music 31 31 31 100.00% 100.00% 0.00%
Note 71 64 66 90.14% 92.96% 3.13%
Photo 18 18 18 100.00% 100.00% 0.00%
Translation 51 48 48 94.12% 94.12% 0.00%
Shopping 38 36 36 94.74% 94.74% 0.00%
Login 42 40 41 95.24% 97.62% 2.50%
Calculation 57 57 57 100.00% 100.00% 0.00%
Weather 28 27 27 96.43% 96.43% 0.00%
Alarm-clock 33 33 33 100.00% 100.00% 0.00%
Sum/Avg 408 392 396 96.08% 97.06% 1.02%
Across the ten evaluated scenarios, the initial decision accuracy ranges from 90.14% to 100.00%, while the final accuracy ranges from 92.96% to 100%. On average, the initial accuracy ($Acc_{1}$) is 96.82%, and the final accuracy ($Acc_{f}$) increases to 97.59%, demonstrating that the self-correction mechanism yields a measurable improvement in decision reliability. In four scenarios, the Decider achieves perfect accuracy without correction, whereas in three other scenarios, accuracy improves noticeably after correction. This result confirms that the self-correction mechanism effectively compensates for transient decision errors caused by dynamic GUI updates or ambiguous widget states.
4.6. RQ5: Scenario-based Test Generation Efficiency
4.6.1. Token Consumption Analysis
Table 6. Result of RQ5: Token Overhead of Scenario-based Test Generation
Scenario $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$ $T_{total}$
Email 5586 8295 5281 13970 11823 833 45788
Music 3748 10197 4982 14729 8841 1963 44460
Note 6461 13639 6602 21667 12178 2573 63120
Photo 1486 3057 1791 6533 3571 746 17184
Translation 4645 6452 5640 11758 10296 359 39150
Shopping 8782 14461 12845 26102 19164 2757 84111
Login 8434 13263 7830 21329 18205 1253 70314
Calculation 5167 9184 6116 15603 12117 1940 50127
Weather 4522 11104 5474 12961 9419 1696 45176
Alarm-clock 3490 3794 3229 9120 8270 456 28359
Average 5232 9345 5979 15377 11388 1458 48779
4.6.2. Time Overhead Analysis
Table 7. Result of RQ5: Time Overhead of Scenario-based Test Generation (minute)
Scenario GPTDroid ScenTest ASOT ScenGen
Email 6.77 1.27 8.46 4.22
Music 7.13 1.24 7.63 3.37
Note 7.21 1.94 12.37 4.65
Photo 8.97 2.95 13.26 5.86
Translation 7.37 1.35 8.16 4.65
Shopping 6.58 2.16 11.69 4.81
Login 8.71 1.82 12.70 5.55
Calculation 7.36 2.10 10.68 3.15
Weather 6.78 1.30 11.54 5.50
Alarm-clock 7.93 1.05 7.35 5.39
Average 7.48 1.72 10.38 4.71
As shown in Table 7, the average time overhead of ScenGen (4.71 minutes) is higher than that of ScenTest (1.72 minutes) but significantly lower than that of ASOT (10.38 minutes) and GPTDroid (7.48 minutes). This demonstrates that ScenGen achieves a balanced trade-off between test generation efficiency and scenario reasoning capability. While its runtime exceeds ScenTest due to additional LLM interactions and vision-language inference, the increased computational effort translates directly into superior coverage and success rates, as demonstrated in RQ1.
RQ5 confirms that ScenGen achieves efficient and resource-aware scenario-based test generation. Its token consumption scales proportionally with scenario complexity, reflecting effective resource allocation across different reasoning stages. In terms of time overhead, ScenGen demonstrates moderate runtime with stable performance across diverse app scenarios, significantly faster than other advanced LLM-based frameworks while maintaining superior reliability and adaptability. These findings affirm that ScenGen provides a computationally efficient yet semantically powerful solution for real-world scenario-based GUI testing.
@@ 6.1. Vision-based GUI Testing
Early efforts primarily focus on visual defect detection and widget recognition. Liu et al. (Liu et al., 2020) propose OwlEye, a tool that detects display inconsistencies in GUI screenshots. White et al. (White et al., 2019) apply machine learning techniques to identify GUI widgets, thereby improving the efficiency of random testing strategies. Xie et al. (Xie et al., 2020) introduce UIED, which integrates traditional computer vision with deep learning to enhance widget detection accuracy, while Zhang et al. (Zhang et al., 2021) develop a more efficient model for GUI element detection. Mansur et al. (Hasan Mansur et al., 2023) present AidUI, which combines CV and NLP techniques to detect dark patterns, improving the ethical and usability evaluation of user interfaces. Besides, Baral et al. (Baral et al., 2024) automate the GUI-based test oracle construction for mobile apps, further illustrating the potential of vision-driven approaches in improving testing reliability.
@@ Acknowledgements.
The authors would like to thank the editors and anonymous reviewers for their time and comments. This work is partially supported by the National Key Research and Development Program of China (2024YFF0908005).